OpenStudy (anonymous):

Please read my replies after as well: (a) Suppose on Jan 1, 1997 Dave invested $2,000 into a bank account at 5% interest compounded continuously. Let y(t) be the value of Dave's investment after t years. Give an exact formula for y(t) (b) Also on Jan 1, 1997 John decides to invest. He put $2,500 into an account at 3% interest compounded monthly. Let g(t) be the value of John's investment after t years. Give an exact formula for g(t). (c) Which account is worth more after 9 years? [ Must show work to receive credit.] (d) To the nearest tenth, at what time t is the value of both accounts the same?

OpenStudy (anonymous):

I have the answers too everything except D, i dont understand why i cant get it

OpenStudy (anonymous):

a) y(t) = 2000e^.05t b) g(t) = 2500(1+.03/12)^12t c) g(t) is larger d) ???

OpenStudy (anonymous):

i set the 2 equations equal to each other, but i cannot isolate t... not sure if im doing it wrong or idk lol help?

OpenStudy (anonymous):

what do you mean? that would give me the answer to c, 3136.62 and 3273.81

OpenStudy (anonymous):

yeah, thats the one i can not get any thoughts?

OpenStudy (anonymous):

what i did was 2000e^.05t = 2500 (1+.03/12)^12(t) i set them equal to each other because if im trying to find when they both give the same answer, that means they themselves are equal.. but i cant find a way to isolate T in order to find how long it takes

OpenStudy (anonymous):

yeah i got to that part too... took the ln of both sides and got that, but still can't isolate the t to find what it's worth. where do i go from there?

OpenStudy (anonymous):

wait what happened to the 1.0025?

OpenStudy (anonymous):

sorry i still dont see how that brings me to the answer, the answer rounded should come out to 11.1

OpenStudy (anonymous):

lol alright let me know what you come up with

OpenStudy (anonymous):

anything?

OpenStudy (anonymous):

sigh =(.. this is ridiculous. but im right in setting the equations equal to each other right?

OpenStudy (anonymous):

You're absolutely right!

OpenStudy (anonymous):

this is really frustrating... thank you for your help tho, i hope its something wrong with the problem lol

OpenStudy (anonymous):

thank you, sorry for having you spend all that time on me

OpenStudy (anonymous):

e^.05t = 1.25 ( 1.0025 )^ 12t .05t = ln 1.25 + 12 ln ( 1.0025 ) t .[ 05 - 12 ln ( 1.0025) ] t = ln 1.25 => t = 11.136

OpenStudy (anonymous):

Finally, I found my flaw =)

OpenStudy (anonymous):

how did the addition sign come along?

OpenStudy (anonymous):

ln ( xy) = ln x + lny

OpenStudy (anonymous):

oh wow dude you're a boss. how did i not think of that -__-... thanks a lot man haha much appreciated

OpenStudy (anonymous):

lol i could imagine its 3 am and i have this final tomorrow at 4 haha

OpenStudy (anonymous):

Rutgers university?

OpenStudy (anonymous):

or do you mean time wise? lol

OpenStudy (anonymous):

oh haha yeah jersey strong =P

OpenStudy (anonymous):

i actually do have a problem i dont understand, i need to figure out how to show you the graph tho, let me see if i can save the picture

OpenStudy (anonymous):

Open the new post to attract more help :)

OpenStudy (anonymous):

lol alright

OpenStudy (anonymous):

woop, thats the entire exam review, its number 25 lol